# Benchmarking Vision Foundation Models for Input Monitoring in Autonomous Driving

### v2

@@ Abstract
Deep neural networks (DNNs) remain challenged by distribution shifts in complex open-world domains like automated driving (AD): Absolute robustness against yet unknown novel objects (semantic shift) or styles like lighting conditions (covariate shift) cannot be guaranteed. Hence, reliable operation-time monitors for identification of out-of-training-data-distribution (OOD) scenarios are imperative. Current approaches for OOD classification are untested for complex domains like AD, are limited in the kinds of shifts they detect, or even require supervision with OOD samples. To prepare for unanticipated shifts, we instead establish a framework around a principled, unsupervised and model-agnostic method that unifies detection of all kinds of shifts: Find a full model of the training data’s feature distribution, to then use its density at new points as in-distribution (ID) score. To implement this, we propose to combine the newly available Vision Foundation Models (VFM) as feature extractors with one of four alternative density modeling techniques. In an extensive benchmark of 4 VFMs against 20 baselines, we show the superior performance of VFM features encodings compared to shift-specific OOD monitors. Additionally, we find that sophisticated architectures outperform larger latent space dimensionality; and our method identifies samples with higher risk of errors on downstream tasks, despite being model-agnostic. This suggests that VFMs are promising to realize model-agnostic, unsupervised, reliable safety monitors in complex vision tasks.
@@ 1 Introduction
Figure 1: A monitoring system for autonomous driving that uses a pre-trained Vision Foundation Model’s image encoder to detect input anomalies that differ from the network’s training data.
@@
In recent decades, deep neural networks (DNNs) have become widely used for environmental perception due to their superior performance [11, 14, 89, 56, 72, 47, 44]. However, in complex open-world domains such as AD and robot navigation this performance is inevitably challenged by the “long tail” of the input distribution, i.e., unforeseeable OOD scenarios for which DNN generalization cannot be guaranteed. To still ensure safe system functioning, reliable monitors that can trigger corrective system measures like handover have become mandatory [16]. Aside from model uncertainty and behavior monitoring [60, 24], an important model-agnostic approach for OOD detection is input monitoring (Fig. 1). At its core it detects when input data fed to the system deviates from the training data distribution [60]. Monitoring methods include training models to recognize historical successes and failures, issuing alerts for failure cases [67, 63], and tracking deviations in the input domain to flag abnormal samples [28].
Unfortunately, existing methods are often highly specialized on the model type, the type of shift, or even specific examples of shifts supplied during monitor training. This sacrifices reliability for completely unknown distribution shifts. A principled way to capture these would be to fully model the training image feature distribution, such that the density value respectively likelihood of a new samples can serve as its ID-score. For complex input domains, however, this requires a powerful feature extractor; and standard DNNs or autoencoders (AE) trained on a dataset of limited diversity fail to provide this. This is the point, where the rich feature encodings of the newly available Vision Foundation Models (VFM) might come to the rescue.
Foundation Models (FM) such as DINO [8] are widely understood as large-scale machine learning models trained on vast amounts of diverse data. They possess the ability to be applied to various downstream tasks, not necessarily directly related to the original training objective [83]. Special appeal of VFMs for input monitoring in AD systems arises from their strong performance across a wide range of visual tasks and their ability to generalize to novel scenarios [34, 52].
In this study we investigate: How suitable are VFM image encodings for modeling a data distribution for OOD scenario detection in complex vision domains like AD?
@@
In particular: How do they compare to standard DNN encodings and shift-specific baseline OOD detection methods? Is this OOD of use for improving safety in downstream applications? Also, while most existing works focuses on either semantic OOD detection on object level [9, 5, 10] or on covariate shifts detection [6, 30], we are interested in monitoring capabilities for both semantic, covariate, and combined shift detection on scene level. Our concrete contributions to this are:
@@
    We propose a modular pipeline for unsupervised, model-agnostic detection of any distribution shift in input images (see Fig. 1), in particular both semantic, covariate, and joint shifts. At its core we use rich VFM feature encodings for distribution density modeling.
@@
    We for the first time benchmark VFMs for OOD on the challenging open-world domain of AD.
    Our extensive tests compare 4 VFMs on 9 both synthetic and real datasets, and 7 shift-specific OOD methods as baselines.
@@
    Our results show the clear superiority of VFM encodings in unified distribution shift detection. In particular, sophisticated architectures consistently outperform larger latent space dimensionality for all shift types.
@@
    We demonstrate practical usability on a downstream AD perception application, showcasing that—despite being model-agnostic—VFMs successfully identify OOD samples with higher risk of errors.
  5. (v)
    We establish a comprehensive benchmarking framework for evaluating input monitoring methods in AD by providing diverse experimental scenarios combining different types of distribution shifts, datasets, and evaluation protocols.
Figure 2: An example of three types of data that can organically arise when deploying models in the open world: (1) ID data (e.g., traffic scene in daylight), (2) covariate shift (e.g., traffic scene under the rain), and (3) semantic shift data (e.g., a tiger on the road).
@@ 2 Related Work
Semantic Shift in Autonomous Driving
In machine learning, training and testing sets are typically drawn from the same distribution, which essentially poses the closed-world assumption [60] that the set of considered semantic classes will not change. [57, 31]. To still identify semantic shifts, i.e., novel classes, techniques like Monte Carlo (MC) dropout estimate neural network uncertainty, eliminating the need for separate OOD test sets [20]. It is used in applications such as semantic segmentation [38] and Mask-RCNN for OOD scenarios [26, 29]. Additionally, advancements include integrating semantic segmentation with anomaly detection [5, 10], and employing generative and reconstruction methods like JSR-Net [80], autoencoder-based approaches [53], and structure-enhanced recurrent VAEs [82] to enhance OOD detection on object level.
Feature extraction techniques are also utilized to identify OOD samples. SAFE [81] teaches an object detector to recognize sensitivity-aware features by perturbing the input dataset. A surrogate model then learns the features of ID data with perturbations, eliminating the need for OOD data. Lastly, FOOD-ID [41] employs a likelihood-based mechanism on the feature space of the classes to identify OOD samples. However, current solutions to semantic shift detection often rely on specific architectures, OOD training data, computationally expensive generative models, or are specific to certain (anticipatable!) distribution shift types. By leveraging pre-trained, off-the-shelf models, scenes with distribution shifts can be detected in an unsupervised and model-agnostic manner.
Covariate Shift in Autonomous Driving
The distribution of real-world scenarios styles diverges from the training data distribution, a phenomenon known as covariate shift [68], also called distributional shift [30] or domain shift [48]. In autonomous driving, covariate shift occurs, e.g., due to changes in lighting (day vs. night) and weather (rain, fog, snow). Several studies have investigated this in AD datasets. The comparison in [30] found that likelihood regret [84] and the self-supervision from [70] outperform variational autoencoders [74] for covariate shift tasks. Lohdefink et al. [48] used a self-supervised approach with Earth Mover’s Distance to link domain mismatch with performance degradation. Bolte et al. [6] applied mean squared error on feature maps to detect covariate shift, while Zhang et al. [87] presented DeepRoad, an unsupervised validation framework for AD systems using generative adversarial networks and metamorphic testing. Just as for semantic shift, current solutions often are model-specific and/or -shift, supervised, or expensively relying on generative models. While VFM inference might still be expensive, we aim to tackle the other limitations in a first step in this work.
Vision Foundation Models
Foundation models (FMs) are large-scale machine learning models trained on diverse data, allowing them to adapt to varied tasks, even those beyond their original objective [83]. The core model, often trained with self-supervised learning [27, 13] or contrastively with (image,caption) pairs [59], can be fine-tuned for tasks like counting novel objects [2], text-based image retrieval [75, 73], and zero-shot classification [59]. In this work, we used the image encoders of FMs like CLIP [59], DINO [8], DINOv2 [54], and Grounding-DINO [44]. CLIP[59] is a well-known VFM trained contrastively on large image-text datasets [69]. Through cosine distance minimization in a shared embedding space, visual and textual representations are aligned in CLIP, enabling zero-shot transfer and robust semantic understanding across diverse vision tasks. DINO [8] uses a self-supervised, label-free learning framework with self-distillation, producing highly semantic features well-suited for transfer learning. Building on this, DINOv2 [54] enhances stability and training speed, supports larger models and datasets, and incorporates masked image modeling, resulting in stronger and more transferable visual representations. Grounding DINO [44] extends DINO [85] with a text-guided object detection framework, enabling precise object localization based on text prompts.
@@ 3 Shift Detection Methodology & Evaluation
Shift Detection Methodology: Overview
@@
We consider the following choices for the monitor distribution modeling further detailed in the following in order of increasing complexity: 1) instance-based: averaged pairwise similarities to the training examples, 2) a one-class support vector machine (OC-SVM), 3) a Gaussian mixture model (GMM) and 4) a normalizing flow (NF).
Average Pairwise Similarity (APS): In this method, we compute the similarity between a test image and training images using cosine similarity. The test set includes both $\mathcal{X}_{\mathrm{ID}}$ and $\mathcal{X}_{\mathrm{OOD}}$ images. Cosine similarity is chosen for several reasons: (1) its effectiveness in measuring the similarity between high-dimensional vectors, particularly when the magnitude of the vectors is less important than their direction; (2) its invariance to scaling, which is beneficial when dealing with normalized feature vectors; and (3) its natural alignment with the latent space of many pre-trained foundational models used for feature extraction. Notably, models like CLIP and DINO are often trained using contrastive losses based on cosine similarity, making this metric particularly suitable for comparing features in their embedding spaces. For a new test image $x^{*}$ with feature vector $f^{*}$, we calculate its average cosine similarity to all training samples
@@
One-Class Support Vector Machine (OC-SVM): OC-SVM [7, 1] is an unsupervised machine learning algorithm designed for novelty detection. It is particularly effective in scenarios where there is an abundance of data from one class (the ”normal” class) and the objective is to identify instances that deviate from this norm. In our context, we employ OC-SVM to model the distribution of features extracted from $\mathcal{X}_{\mathrm{monitor}}$ images and subsequently detect samples from $\mathcal{X}_{\mathrm{ID}}$ and $\mathcal{X}_{\mathrm{OOD}}$ sets. For a set of $\mathcal{X}_{\mathrm{monitor}}$ features $\mathcal{F}_{\mathrm{monitor}}=\{f_{1},f_{2},\ldots,f_{n}\}$ extracted from $\mathcal{X}_{\mathrm{monitor}}$ set, the OC-SVM aims to find a hyperplane that separates the data from the origin with maximum margin in the feature space. This is achieved by solving the following optimization problem:
$\min_{w,\xi,\rho}\frac{1}{2}||w||^{2}+\frac{1}{\nu n}\sum_{i=1}^{n}\xi_{i}-\rho$
subject to: $(w\cdot\phi(f_{i}))\geq\rho-\xi_{i},\quad\xi_{i}\geq 0$
@@
$\SwapAboveDisplaySkip c_{\text{OC-SVM}}(x^{*})=(w\cdot\phi(f^{*}))-\rho$
@@
Gaussian Mixture Model (GMM): We model the features $\mathcal{F}_{\mathrm{train}}$ using a Gaussian mixture model (GMM) of $K$ Gaussians with the formulation
@@
The parameters $\Theta$ of the GMM are derived from the feature set $\mathcal{F}_{\mathrm{train}}$ using the Expectation-Maximization (EM) algorithm. To optimally balance complexity and fitting accuracy, the model’s number of components $K$ is determined through the Akaike Information Criterion (AIC). This criterion is applied by testing a series of $K$ values on $\mathcal{F}_{\mathrm{monitor}}$. The choice of $K$ is discussed in the supplementary material.
@@
$\SwapAboveDisplaySkip c_{\text{GMM}}(x^{*})=\log p(f^{*}|\Theta)=\log\left(\sum_{k=1}^{K}\pi_{k}\mathcal{N}(f^{*}|\mu_{k},\Sigma_{k})\right)$
@@
Normalizing Flows (NF): Normalizing flows [55, 18] provide a deep learning framework for high-dimensional density estimation through invertible mappings. Let $P_{F}$ denote the probability measure of VFM features $f_{1},\ldots,f_{n}$ from $\mathcal{F}_{\mathrm{monitor}}$, and $P_{\mathcal{N}}$ denote the measure of the multivariate standard normal distribution $\mathcal{N}(\cdot|0,I)$. A normalizing flow learns an invertible map $g_{\theta}$ such that:

### v3

@@ Abstract
Deep neural networks (DNNs) remain challenged by distribution shifts in complex open-world domains like automated driving (AD): Robustness against yet unknown novel objects (semantic shift) or styles like lighting conditions (covariate shift) cannot be guaranteed. Hence, reliable operation-time monitors for identification of out-of-training-data-distribution (OOD) scenarios are imperative. Current approaches for OOD classification are untested for complex domains like AD, are limited in the kinds of shifts they detect, or even require supervision with OOD samples. To prepare for unanticipated shifts, we instead establish a framework around a principled, unsupervised and model-agnostic method that unifies detection of semantic and covariate shifts: Find a full model of the training data’s feature distribution, to then use its density at new points as in-distribution (ID) score. To implement this, we propose to combine Vision Foundation Models (VFMs) as feature extractors with density modeling techniques. Through a comprehensive benchmark of 4 VFMs with different backbone architectures and 5 density-modeling techniques against established baselines, we provide the first systematic evaluation of OOD classification capabilities of VFMs across diverse conditions. A comparison with state-of-the-art binary OOD classification methods reveals that VFM embeddings with density estimation outperform existing approaches in identifying OOD inputs. Additionally, we show that our method detects high-risk inputs likely to cause errors in downstream tasks, thereby improving overall performance. Overall, VFMs, when coupled with robust density modeling techniques, are promising to realize model-agnostic, unsupervised, reliable safety monitors in complex vision tasks.
@@ 1 Introduction
Figure 1: A monitoring system for autonomous driving that uses a pre-trained Vision Foundation Model’s image encoder to detect distribution shifts in input data. The system identifies whether camera inputs (red dots) fall within the model’s training distribution (brown dots) or represent covariate shifts (top) or semantic shifts (bottom), helping assess operational safety and potential failure risks.
@@
In recent decades, deep neural networks (DNNs) have become widely used for environmental perception due to their superior performance compared to traditional computer vision methods [13, 15, 90, 58, 73, 48, 45]. However, in complex open-world domains such as autonomous driving (AD) and robot navigation this performance is inevitably challenged by the “long tail” of the input distribution, i.e., unforeseeable out-of-distribution (OOD) scenarios for which DNN generalization cannot be guaranteed. AD spans multiple operational levels, from driver assistance (L1-L2) to fully autonomous operation (L4-L5), with conditional automation (L3) bridging the transition [67]. Across these levels, camera-based perception systems support critical functions, from lane-keeping and traffic sign recognition to full scene understanding [9]. As these systems operate in increasingly diverse environments, a fundamental question emerges: How can we ensure that perception models function reliably in complex, open-world scenarios, particularly when encountering inputs that deviate from their training distribution? To mitigate these risks and ensure safe operation, reliable monitors that can detect and respond to such deviations have become mandatory [17, 34]. Aside from model uncertainty and behavior monitoring [62, 25], an important approach for OOD detection is input monitoring (Fig. 1), which identifies when input data fed to the system deviates from the training data distribution [62].
Existing binary OOD classification methods suffer from critical limitations for AD applications: they often require direct access to model internals, are specialized to specific architectures, or need exposure to anticipated shift types during training [32, 46, 51, 89, 76]. These constraints compromise reliability when encountering truly unforeseen distribution shifts in open-world environments – a critical limitation for autonomous driving systems that must handle diverse operational domains. Furthermore, accessing internal model layers becomes impractical in AD systems. A more principled approach would model the complete training data distribution in feature space, using density values or likelihoods as in-distribution (ID) scores for new samples. The challenge for complex visual domains lies in extracting sufficiently powerful features—standard DNNs and autoencoders (AE) trained on limited datasets typically fail to capture the necessary representational richness.
This is precisely where VFMs offer a compelling solution. Trained on diverse, large-scale datasets, VFMs encode comprehensive semantic contexts and subtle visual distinctions within their latent representations [10, 56], making them particularly well-suited for distribution modeling. Their appeal for input monitoring in AD stems from demonstrated strong generalization capabilities across diverse visual tasks and novel scenarios [33, 54], potentially enabling robust detection of unforeseen distribution shifts without requiring model-specific adaptations or prior exposure to OOD examples.
@@
In this study we investigate: How suitable are VFM image encodings for modeling a data distribution for OOD scenario detection in complex vision domains like AD? In particular: How do they compare to standard DNN encodings and SoTA baseline binary-OOD detection methods? Is this OOD of use for improving overall system reliability in downstream applications? Also, while most existing works focuses on either semantic OOD detection on object level [11, 5, 12] or on covariate shifts detection [6, 29], we are interested in monitoring capabilities for both semantic, covariate, and combined shift detection on scene level. Our contributions are:
@@
    We propose a modular pipeline for unsupervised, model-agnostic detection of distribution shifts in input images (see Fig. 1), in particular both semantic, covariate, and joint shifts. At its core we use rich VFM feature encodings for distribution density modeling.
@@
    We for the first time benchmark VFMs for OOD on the challenging open-world domain of AD. Our extensive tests compare 4 VFMs on 12 both synthetic and real datasets, and 7 OOD methods as baselines. Our results show the clear superiority of VFM encodings in unified distribution shift detection.
@@
    We demonstrate practical usability on a downstream AD perception application, showcasing that, despite being model-agnostic, VFMs successfully identify OOD samples with higher risk of errors.
@@
    We establish a comprehensive benchmarking framework for evaluating binary OOD classification methods in AD by providing diverse experimental scenarios combining different types of distribution shifts, datasets, and evaluation protocols.
@@ 2 Related Work
2.1 Semantic Shift in Autonomous Driving
In machine learning, training and testing sets are typically drawn from the same distribution, which poses the closed-world assumption [62] that the set of considered semantic classes will not change [59, 30]. To identify semantic shifts (novel classes), techniques like Monte Carlo dropout estimate neural network uncertainty [21], finding applications in semantic segmentation [38] and object detection [27, 28]. Recent advances integrate semantic segmentation with anomaly detection [5, 12] and employ various reconstruction approaches [79, 55, 82] for object-level OOD detection. Alternative feature-based approaches include SAFE [81], which teaches detectors to recognize sensitivity-aware features through input perturbation, and FOOD-ID [41], which employs likelihood-based feature space mechanisms. However, existing solutions typically rely on model-specific architectures, OOD training examples, computationally expensive generative models, or are limited to specific anticipatable shift types. In contrast, pre-trained VFMs enable unsupervised, model-agnostic detection of distribution shifts without these limitations.
2.2 Covariate Shift in Autonomous Driving
The distribution of real-world scenarios often diverges from training data, a phenomenon known as covariate shift [70], distributional shift [29], or domain shift [49]. In autonomous driving, this occurs due to changes in lighting conditions and weather patterns. Comparative studies [29] found likelihood regret [85] and self-supervised approaches [71] outperform variational autoencoders [74] for covariate shift detection. Other methods include Earth Mover’s Distance for linking domain mismatch with performance degradation [49], feature map error analysis [6], and GAN-based validation frameworks [88].
Similar to semantic shift detection, these approaches rely on model-specific architectures, require supervision with specific shift examples, or depend on computationally expensive generative models. Our work addresses these fundamental limitations by leveraging VFMs’ rich representational power for model-agnostic distribution shift detection that unifies both covariate and semantic shift identification.
2.3 Vision Foundation Models
FMs are large-scale machine learning models trained on diverse data, allowing them to adapt to varied tasks beyond their original objective [83]. The field includes models like CLIP [61], trained contrastively on image-text pairs; DINO [10], using self-supervised learning with self-distillation; DINOv2 [56], enhancing DINO with masked image modeling; Grounding-DINO [45], extending DETR-based detection [86] with text guidance; InternImage [80], utilizing deformable convolutions for adaptive receptive fields; Florence-2 [84], employing hierarchical visual tokenization; and LLaVA [44], connecting vision encoders to large language models through visual instruction tuning. This study focuses on the first four VFMs, selected for their architectural diversity (spanning Convolutional Networks, Vision Transformers, and Swin Transformers) and absence of test data leakage (see Supplementary). For a comprehensive overview of FMs, we refer to [7, 2], while [91] provides a detailed review of VFMs specifically in AD.
@@ 3 Shift Detection Methodology & Evaluation
3.1 Shift Detection Methodology: Overview
@@
We consider the following choices for the monitor distribution modeling further detailed in the following in order of increasing complexity: 1) instance-based: averaged pairwise similarities to the training examples and its computationally efficient derivation using similarity to the mean feature vector, 2) a one-class support vector machine, 3) a Gaussian mixture model and 4) a normalizing flow.
Average Pairwise Similarity (APS) and Mean Feature Similarity (MFS):
In this method, we compute the similarity between a test image and training images using cosine similarity. The test set includes both $\mathcal{X}_{\mathrm{ID}}$ and $\mathcal{X}_{\mathrm{OOD}}$ images. Cosine similarity is chosen for several reasons: (1) its effectiveness in measuring the similarity between high-dimensional vectors, particularly when the magnitude of the vectors is less important than their direction; (2) its invariance to scaling, which is beneficial when dealing with normalized feature vectors; and (3) its natural alignment with the latent space of many pre-trained foundational models used for feature extraction. Notably, models like CLIP and DINO are often trained using contrastive losses based on cosine similarity, making this metric particularly suitable for comparing features in their embedding spaces. For a new test image $x^{*}$ with feature vector $f^{*}$, we calculate its average cosine similarity to all training samples.
@@
For real-time applications, we also implement MFS, a computationally efficient variant that first computes a mean feature vector $\bar{f}=\frac{1}{n}\sum_{i=1}^{n}f_{i}$ from training samples, then calculates similarity to this prototype:
$c_{\text{MFS}}(x^{*})=\cos(f^{*},\bar{f})=\frac{f^{*}\cdot\bar{f}}{||f^{*}||\,||\bar{f}||}$ (2)
This reduces inference complexity from $O(n)$ to $O(1)$, making it suitable for monitoring systems with constrained computational resources.
One-Class Support Vector Machine (OC-SVM):
OC-SVM [8, 1] is an unsupervised machine learning algorithm designed for novelty detection. It is particularly effective in scenarios where there is an abundance of data from one class (the ”normal” class) and the objective is to identify instances that deviate from this norm. In our context, we employ OC-SVM to model the distribution of features extracted from $\mathcal{X}_{\mathrm{monitor}}$ images and subsequently detect samples from $\mathcal{X}_{\mathrm{ID}}$ and $\mathcal{X}_{\mathrm{OOD}}$ sets. For a set of $\mathcal{X}_{\mathrm{monitor}}$ features $\mathcal{F}_{\mathrm{monitor}}=\{f_{1},f_{2},\ldots,f_{n}\}$ extracted from $\mathcal{X}_{\mathrm{monitor}}$ set, the OC-SVM aims to find a hyperplane that separates the data from the origin with maximum margin in the feature space. This is achieved by solving the following optimization problem:
$\min_{w,\xi,\rho}\frac{1}{2}||w||^{2}+\frac{1}{\nu n}\sum_{i=1}^{n}\xi_{i}-\rho$ (3)
subject to: $(w\cdot\phi(f_{i}))\geq\rho-\xi_{i},\quad\xi_{i}\geq 0$ (4)
@@
$\SwapAboveDisplaySkip c_{\text{OC-SVM}}(x^{*})=(w\cdot\phi(f^{*}))-\rho$ (5)
@@
Gaussian Mixture Model (GMM):
We model the features $\mathcal{F}_{\mathrm{train}}$ using a Gaussian mixture model (GMM) [52] of $K$ Gaussians with the formulation
@@
The parameters $\Theta$ of the GMM are derived from the feature set $\mathcal{F}_{\mathrm{train}}$ using the Expectation-Maximization (EM) algorithm. To optimally balance complexity and fitting accuracy, the model’s number of components $K$ is determined through the Akaike Information Criterion (AIC). This criterion is applied by testing a series of $K$ values on $\mathcal{F}_{\mathrm{monitor}}$. The choice of $K$ is discussed in the Appendix.
@@
$\SwapAboveDisplaySkip c_{\text{GMM}}(x^{*})=\log p(f^{*}|\Theta)=\log\left(\sum_{k=1}^{K}\pi_{k}\mathcal{N}(f^{*}|\mu_{k},\Sigma_{k})\right)$ (7)
@@
Normalizing Flows (NF):
Normalizing flows [57, 19] provide a deep learning framework for high-dimensional density estimation through invertible mappings. Let $P_{F}$ denote the probability measure of VFM features $f_{1},\ldots,f_{n}$ from $\mathcal{F}_{\mathrm{monitor}}$, and $P_{\mathcal{N}}$ denote the measure of the multivariate standard normal distribution $\mathcal{N}(\cdot|0,I)$. A normalizing flow learns an invertible map $g_{\theta}$ such that: